# Defensive Unlearning with Adversarial Training for Robust Concept Erasure in Diffusion Models

### v2

@@ Abstract
Diffusion models (DMs) have achieved remarkable success in text-to-image generation, but they also pose safety risks, such as the potential generation of harmful content and copyright violations. The techniques of machine unlearning, also known as concept erasing, have been developed to address these risks. However, these techniques remain vulnerable to adversarial prompt attacks, which can prompt DMs post-unlearning to regenerate undesired images containing concepts (such as nudity) meant to be erased. This work aims to enhance the robustness of concept erasing by integrating the principle of adversarial training (AT) into machine unlearning, resulting in the robust unlearning framework referred to as AdvUnlearn. However, achieving this effectively and efficiently is highly non-trivial. First, we find that a straightforward implementation of AT compromises DMs’ image generation quality post-unlearning. To address this, we develop a utility-retaining regularization on an additional retain set, optimizing the trade-off between concept erasure robustness and model utility in AdvUnlearn. Moreover, we identify the text encoder as a more suitable module for robustification compared to UNet, ensuring unlearning effectiveness. And the acquired text encoder can serve as a plug-and-play robust unlearner for various DM types. Empirically, we perform extensive experiments to demonstrate the robustness advantage of AdvUnlearn across various DM unlearning scenarios, including the erasure of nudity, objects, and style concepts. In addition to robustness, AdvUnlearn also achieves a balanced tradeoff with model utility. To our knowledge, this is the first work to systematically explore robust DM unlearning through AT, setting it apart from existing methods that overlook robustness in concept erasing. Codes are available at: https://github.com/OPTML-Group/AdvUnlearn
@@ 1 Introduction
Recent rapid advancements in diffusion models (DMs) (Ho et al., 2020; Song and Ermon, 2019; Song et al., 2020; Dhariwal and Nichol, 2021; Nichol and Dhariwal, 2021; Watson et al., 2021; Rombach et al., 2022; Croitoru et al., 2023) have popularized the realm of text-to-image generation. These models, trained on extensive online datasets, can generate remarkably realistic images. However, their training heavily relies on diverse internet-sourced content and can introduce safety concerns when prompted with inappropriate texts, such as the generation of NSFW (Not Safe For Work) images, highlighted in several studies (Rando et al., 2022; Schramowski et al., 2023). To address this concern, post-hoc safety checkers were initially applied to DMs (Schramowski et al., 2023; Nichol et al., 2021). However, they were later found to be inadequate in effectively preventing the generation of unsafe content. To further enhance safety, the concept of machine unlearning (MU) has been introduced (Nguyen et al., 2022; Shaik et al., 2023; Cao and Yang, 2015; Thudi et al., 2022; Xu et al., 2023; Jia et al., 2023; Fan et al., 2024), aiming to mitigate the influence of undesired textual concepts in DM training or fine-tuning (Gandikota et al., 2023a; Zhang et al., 2023a; Kumari et al., 2023a; Gandikota et al., 2023b). As a result, DMs post-unlearning (referred to as ‘concept-erased DMs’ or ‘unlearned DMs’) are designed to negate the generation of undesirable content, even when faced with inappropriate prompts.
Despite the recent progress with unlearned, safety-driven DMs, recent studies (Zhang et al., 2023b; Chin et al., 2023; Tsai et al., 2023; Pham et al., 2024a) have shown that these models remain vulnerable to generating unsafe images when exposed to adversarial prompt attacks, which involve minor adversarial perturbations in input prompts. These attacks can readily jailbreak concept-erased DMs to regenerate content subject to the concept targeted for unlearning, even if these DMs perform well against inappropriate prompts in a non-adversarial context. In Fig. 1, we exemplify the generation of the stable diffusion (SD) v1.4 model before and post unlearning the ‘nudity’ concept. The unlearned model is confronted with an inappropriate prompt from the I2P dataset (Schramowski et al., 2023) and its adversarial prompt counterpart, generated using the attack method UnlearnDiffAtk (Zhang et al., 2023b). The lack of robustness in concept erasing (or machine unlearning) in DMs gives rise to the key research question tackled in this work:
@@
To address (Q), we take inspiration from the successes of adversarial training (AT) (Madry et al., 2017) in enhancing the adversarial robustness of image classification models. To the best of our knowledge, we are the first to study the integration of AT into DM unlearning systematically and to develop a successful integration scheme, termed AdvUnlearn, by addressing its unique effectiveness and efficiency challenges, such as balancing the preservation of image generation quality and selecting the appropriate module to optimize during AT. In the literature, the most relevant work to ours is (Huang et al., 2023), which employs AT to train robust adapters within UNet for DMs. However, our work significantly differs from (Huang et al., 2023). First, we aim for a comprehensive study of AT for DMs, focusing not only on when AT is (in)effective for DMs and why this (in)effectiveness occurs but also on how to improve it. Additionally, we explore which advancements in AT for robust image classification can be translated into improving the robustness of DMs. Second, we identify that retaining image generation quality is a major challenge when integrating AT into DMs, especially in compatibility with DM unlearning methods. We tackle this challenge by drawing inspiration from the AT principle that ‘unlabeled data improves the robustness-accuracy tradeoff’ (Zhang et al., 2019; Carmon et al., 2019; Alayrac et al., 2019), and accordingly develop a utility-retaining regularization scheme based on an augmented retain prompt set. Third, by dissecting DMs into text encoder and UNet components, we discover that the integration of AT with DM unlearning particularly favors the text encoder module. This contrasts with conventional DM unlearning methods, which are typically applied to the UNet.
@@ 2 Related Work
Machine unlearning for concept erasing in DMs. Generating visually authentic images from textual descriptions remains a compelling challenge in generative AI. DMs (diffusion models) have notably advanced, surpassing generative adversarial networks (GANs) in various aspects, particularly in conditional generation subject to text prompts (Kang et al., 2023; Tao et al., 2023, 2022; Li et al., 2019; Xu et al., 2018; Reed et al., 2016; Zhou et al., 2022; Kawar et al., 2023; Zhang et al., 2022a). Despite their success, DMs also present safety and ethics concerns, especially in generating images with harmful content when conditioned on inappropriate prompts (Rando et al., 2022; Schramowski et al., 2023). To address these concerns, several approaches have been proposed, including post-image filtering (Rando et al., 2022), modifying inference guidance (Schramowski et al., 2023), and retraining with curated datasets (Rombach et al., 2022). However, lightweight interventions like the first two may not fully address the model’s inherent propensity to generate controversial content (Schramowski et al., 2023; Birhane et al., 2021; Somepalli et al., 2023). MU (machine unlearning) (Cao and Yang, 2015; Ginart et al., 2019; Jia et al., 2024) is another emerging approach for ensuring safe image generation by erasing the influence of undesired concepts, also referred to as concept erasing in DMs. Leveraging MU principles, various strategies for designing unlearned DMs have been explored, focusing on refining fine-tuning methods such as those by (Gandikota et al., 2023a; Zhang et al., 2023a; Kumari et al., 2023a; Gandikota et al., 2023b; Fan et al., 2023; Lyu et al., 2023; Huang et al., 2023; Kumari et al., 2023b; Heng and Soh, 2023; Ni et al., 2023; Zhang et al., 2024a; Wu et al., 2024; Wu and Harandi, 2024; Pham et al., 2024b). For example, UNets within DMs have been fine-tuned to redirect outputs towards either random or anchored outputs, effectively preventing the generation of images associated with the concepts designated for unlearning (Gandikota et al., 2023a; Zhang et al., 2023a; Kumari et al., 2023a). Additional efforts (Fan et al., 2023; Wu and Harandi, 2024) have employed gradient-based techniques to map out the weight saliency within UNet related to the concept to be unlearned, concentrating fine-tuning efforts on these salient weights. To enhance the efficiency of unlearning, UCE (Gandikota et al., 2023b) introduces a method of closed-form parameter editing specifically for DM unlearning. However, this approach lacks robustness against jailbreaking attacks (Zhang et al., 2023b).
@@
Adversarial prompt attacks against safety-driven DMs. Adversarial prompts or jailbreaking attacks specifically manipulate the text inputs of DMs to produce undesirable outcomes. Similar to the text-based attacks in natural language processing (NLP), adversarial prompt attacks can involve character or word-level manipulations, such as deletions, additions, and substitutions (Eger and Benz, 2020; Liu et al., 2022; Hou et al., 2022; Li et al., 2018; Alzantot et al., 2018; Jin et al., 2020; Garg and Ramakrishnan, 2020; Qiu et al., 2022). Strategies discussed in (Yang et al., 2023) are designed to bypass NSFW safety protocols, cleverly evading content moderation algorithms. Other related attacks (Zhang et al., 2023b; Chin et al., 2023; Tsai et al., 2023; Pham et al., 2024a; Maus et al., 2023; Zhuang et al., 2023) coerce DMs into generating images that contradict their programmed intent. For instance, (Pham et al., 2024a) used textual inversion (Gal et al., 2022) to find a continuous word embedding representing the concept to be unlearned by the model. (Chin et al., 2023) employed ground truth guidance from an auxiliary frozen UNet (Gandikota et al., 2023a) and discrete optimization techniques from (Wen et al., 2024) to craft a white-box adversarial prompt attack. To overcome the dependency on auxiliary model guidance, UnlearnDiffAtk (Zhang et al., 2023b) leveraged the intrinsic classification capabilities of DMs, facilitating the creation of adversarial prompts. In this work, we treat machine unlearning for DMs as a defensive challenge. Our approach involves fine-tuning the target model to not only unlearn specific concepts but also to enhance its robustness against adversarial prompt attacks.
@@
Adversarial training (AT). In the realm of image classification, adversarial attacks that generate subtle perturbations to fool machine learning (ML) models have long posed a robustness challenge for vision systems (Goodfellow et al., 2014; Carlini and Wagner, 2017; Croce and Hein, 2020; Xu et al., 2019; Athalye et al., 2018a; Gong et al., 2022). In response, AT (adversarial training) (Madry et al., 2017), the cornerstone of training-based defenses, conceptualizes defense as a two-player game between the attacker and defender (Madry et al., 2017; Zhang et al., 2019; Carmon et al., 2019; Goodfellow et al., 2014; Xie et al., 2017; Kannan et al., 2018; Athalye et al., 2018b; Stanforth et al., 2019; Cai et al., 2018; Wang et al., 2019). Additionally, TRADES (Zhang et al., 2019) was proposed to strike a better balance between accuracy and robustness. Further studies (Carmon et al., 2019; Alayrac et al., 2019; Stanforth et al., 2019; Zhai et al., 2019) demonstrated that unlabeled data and self-training have proven effective in enhancing robustness and generalization in adversarial contexts. To improve the efficiency of AT, past research also proposed adopting more efficient attack methods or fewer steps to generate adversarial examples (Goodfellow et al., 2014; Wong et al., 2020; Tramèr et al., 2018; Andriushchenko and Flammarion, 2020; Sriramanan et al., 2020, 2021; Zhang et al., 2022b; Tsiligkaridis and Roberts, 2022; Shafahi et al., 2019). In particular, the fast gradient sign method (FGSM) was utilized for adversarial generation in AT (Goodfellow et al., 2014; Wong et al., 2020). And the gradient alignment strategy was proposed to improve the quality of fast AT (Andriushchenko and Flammarion, 2020).
@@ 3 Preliminaries and Problem Statement
Throughout the work, we focus on latent diffusion models (LDMs) (Rombach et al., 2022; Cao et al., 2022), which have exceled in text-to-image generation by integrating text prompts (such as text-based image descriptions) into image embeddings to guide the generation process. In LDMs, the diffusion process initiates with a random noise sample drawn from the standard Gaussian distribution . This sample undergoes a progressive transformation through a series of time steps in a gradual denoising process, ultimately resulting in the creation of a clean image . At each time step , the diffusion model utilizes a noise estimator , parameterized by and conditioned on an input prompt (i.e., associated with a textual concept). The diffusion process operates on the latent representation of the image at each time (). The training objective for is to minimize the denoising error as below:
@@
Concept erasure in DMs. DMs, despite their high capability, may generate unsafe content or disclose sensitive information when given inappropriate text prompts. For example, the I2P dataset (Schramowski et al., 2023) compiles numerous inappropriate prompts capable of leading DMs to generate NSFW content. To mitigate the generation of harmful or sensitive content, a range of studies (Gandikota et al., 2023a; Zhang et al., 2023a; Kumari et al., 2023a; Gandikota et al., 2023b; Fan et al., 2023) have explored the technique of concept erasing or machine unlearning within DMs, aiming to enhance the DM training process by mitigating the impact of undesired textual concepts on image generation.
A widely recognized concept erasing approach is ESD (Gandikota et al., 2023a), notable for its state-of-the-art (SOTA) balance between unlearning effectiveness and model utility preservation (Zhang et al., 2024a). Unless specified otherwise, we will adopt the objective of ESD for implementing concept erasure. ESD facilitates the fine-tuning process of DMs by guiding outputs away from a specific concept targeted for erasure. Let denote the concept to erase, then the diffusion process of ESD is modified to
where denotes the concept-erased DM, is the originally pre-trained DM, and represents unconditional generation of the model by considering text prompt as empty. Compared to the standard conditional DM (Ho and Salimans, 2022) (with classifier-free guidance), the second term encourages the adjustment of the data distribution (with erasing guidance parameter ) to minimize the likelihood of generating an image that could be labeled as . To optimize , ESD performs the following model fine-tuning based on (3):
@@
Adversarial prompts against concept-erased DMs. Although concept erasing enhances safety, recent studies (Zhang et al., 2023b; Chin et al., 2023; Tsai et al., 2023; Pham et al., 2024a) have also shown that concept-erased DMs often lack robustness when confronted with adversarial prompt attacks; see Fig. 1 for examples. Let represent a perturbed text prompt corresponding to , obtained through token manipulation in the text space (Zhang et al., 2023b; Chin et al., 2023) or in the token embedding space (Pham et al., 2024a). The generation of adversarial prompts can be solved as (Zhang et al., 2023b; Chin et al., 2023):
@@
AdvUnlearn: A defensive unlearning setup via AT. The lack of adversarial robustness in concept-erased DMs motivates us to devise a solution that enhances their robustness in the face of adversarial prompts. AT (Madry et al., 2017) offers a principled algorithmic framework for addressing this challenge. It formulates robust concept erasure as a two-player game involving the defender (i.e., the unlearner for concept erasing) and the attacker (i.e., the adversarial prompt). The original AT constrains the attacker’s objective to precisely oppose the defender’s objective. To loosen this constraint, we consider a generalized AT formulation based on bi-level optimization (Zhang et al., 2022b, c, 2023c, 2024b), where the defender and attacker are delineated through the upper-level and lower-level optimization problems, respectively:

### v3

@@ Abstract
Diffusion models (DMs) have achieved remarkable success in text-to-image generation, but they also pose safety risks, such as the potential generation of harmful content and copyright violations. The techniques of machine unlearning, also known as concept erasing, have been developed to address these risks. However, these techniques remain vulnerable to adversarial prompt attacks, which can prompt DMs post-unlearning to regenerate undesired images containing concepts (such as nudity) meant to be erased. This work aims to enhance the robustness of concept erasing by integrating the principle of adversarial training (AT) into machine unlearning, resulting in the robust unlearning framework referred to as AdvUnlearn. However, achieving this effectively and efficiently is highly non-trivial. First, we find that a straightforward implementation of AT compromises DMs’ image generation quality post-unlearning. To address this, we develop a utility-retaining regularization on an additional retain set, optimizing the trade-off between concept erasure robustness and model utility in AdvUnlearn. Moreover, we identify the text encoder as a more suitable module for robustification compared to UNet, ensuring unlearning effectiveness. And the acquired text encoder can serve as a plug-and-play robust unlearner for various DM types. Empirically, we perform extensive experiments to demonstrate the robustness advantage of AdvUnlearn across various DM unlearning scenarios, including the erasure of nudity, objects, and style concepts. In addition to robustness, AdvUnlearn also achieves a balanced tradeoff with model utility. To our knowledge, this is the first work to systematically explore robust DM unlearning through AT, setting it apart from existing methods that overlook robustness in concept erasing. Codes are available at https://github.com/OPTML-Group/AdvUnlearn.
@@ 1 Introduction
Recent rapid advancements in diffusion models (DMs) [1, 2, 3, 4, 5, 6, 7, 8] have popularized the realm of text-to-image generation. These models, trained on extensive online datasets, can generate remarkably realistic images. However, their training heavily relies on diverse internet-sourced content and can introduce safety concerns when prompted with inappropriate texts, such as the generation of NSFW (Not Safe For Work) images, highlighted in several studies [9, 10]. To address this concern, post-hoc safety checkers were initially applied to DMs [10, 11]. However, they were later found to be inadequate in effectively preventing the generation of unsafe content. To further enhance safety, the concept of machine unlearning (MU) has been introduced [12, 13, 14, 15, 16, 17, 18], aiming to mitigate the influence of undesired textual concepts in DM training or fine-tuning [19, 20, 21, 22]. As a result, DMs post-unlearning (referred to as ‘concept-erased DMs’ or ‘unlearned DMs’) are designed to negate the generation of undesirable content, even when faced with inappropriate prompts.
Despite the recent progress with unlearned, safety-driven DMs, recent studies [23, 24, 25, 26] have shown that these models remain vulnerable to generating unsafe images when exposed to adversarial prompt attacks, which involve minor adversarial perturbations in input prompts. These attacks can readily jailbreak concept-erased DMs to regenerate content subject to the concept targeted for unlearning, even if these DMs perform well against inappropriate prompts in a non-adversarial context. In Fig. 1, we exemplify the generation of the stable diffusion (SD) v1.4 model before and post unlearning the ‘nudity’ concept. The unlearned model is confronted with an inappropriate prompt from the I2P dataset [10] and its adversarial prompt counterpart, generated using the attack method UnlearnDiffAtk [23]. The lack of robustness in concept erasing (or machine unlearning) in DMs gives rise to the key research question tackled in this work:
@@
To address (Q), we take inspiration from the successes of adversarial training (AT) [29] in enhancing the adversarial robustness of image classification models. To the best of our knowledge, we are the first to study the integration of AT into DM unlearning systematically and to develop a successful integration scheme, termed AdvUnlearn, by addressing its unique effectiveness and efficiency challenges, such as balancing the preservation of image generation quality and selecting the appropriate module to optimize during AT. In the literature, the most relevant work to ours is [30], which employs AT to train robust adapters within UNet for DMs. However, our work significantly differs from [30]. First, we aim for a comprehensive study of AT for DMs, focusing not only on when AT is (in)effective for DMs and why this (in)effectiveness occurs but also on how to improve it. Additionally, we explore which advancements in AT for robust image classification can be translated into improving the robustness of DMs. Second, we identify that retaining image generation quality is a major challenge when integrating AT into DMs, especially in compatibility with DM unlearning methods. We tackle this challenge by drawing inspiration from the AT principle that ‘unlabeled data improves the robustness-accuracy tradeoff’ [31, 32, 33, 34, 35], and accordingly develop a utility-retaining regularization scheme based on an augmented retain prompt set. Third, by dissecting DMs into text encoder and UNet components, we discover that the integration of AT with DM unlearning particularly favors the text encoder module. This contrasts with conventional DM unlearning methods, which are typically applied to the UNet.
@@ 2 Related Work
Machine unlearning for concept erasing in DMs. Generating visually authentic images from textual descriptions remains a compelling challenge in generative AI. DMs (diffusion models) have notably advanced, surpassing generative adversarial networks (GANs) in various aspects, particularly in conditional generation subject to text prompts [36, 37, 38, 39, 40, 41, 42, 43, 44]. Despite their success, DMs also present safety and ethics concerns, especially in generating images with harmful content when conditioned on inappropriate prompts [9, 10]. To address these concerns, several approaches have been proposed, including post-image filtering [9], modifying inference guidance [10], and retraining with curated datasets [7]. However, lightweight interventions like the first two may not fully address the model’s inherent propensity to generate controversial content [10, 45, 46]. MU (machine unlearning) [14, 47, 48] is another emerging approach for ensuring safe image generation by erasing the influence of undesired concepts, also referred to as concept erasing in DMs. Leveraging MU principles, various strategies for designing unlearned DMs have been explored, focusing on refining fine-tuning methods such as those by [19, 20, 21, 22, 27, 28, 30, 49, 50, 51, 52, 53, 54, 55, 56]. For example, UNets within DMs have been fine-tuned to redirect outputs towards either random or anchored outputs, effectively preventing the generation of images associated with the concepts designated for unlearning [19, 20, 21]. Additional efforts [27, 54] have employed gradient-based techniques to map out the weight saliency within UNet related to the concept to be unlearned, concentrating fine-tuning efforts on these salient weights. To enhance the efficiency of unlearning, UCE [22] introduces a method of closed-form parameter editing specifically for DM unlearning. However, this approach lacks robustness against jailbreaking attacks [23].
@@
Adversarial prompt attacks against safety-driven DMs. Adversarial prompts or jailbreaking attacks specifically manipulate the text inputs of DMs to produce undesirable outcomes. Similar to the text-based attacks in natural language processing (NLP), adversarial prompt attacks can involve character or word-level manipulations, such as deletions, additions, and substitutions [57, 58, 59, 60, 61, 62, 63, 64, 65]. Strategies discussed in [66] are designed to bypass NSFW safety protocols, cleverly evading content moderation algorithms. Other related attacks [23, 24, 25, 26, 67, 68] coerce DMs into generating images that contradict their programmed intent. For instance, Pham et al. [26] used textual inversion [69] to find a continuous word embedding representing the concept to be unlearned by the model. Chin et al. [24] employed ground truth guidance from an auxiliary frozen UNet [19] and discrete optimization techniques from [70] to craft a white-box adversarial prompt attack. To overcome the dependency on auxiliary model guidance, UnlearnDiffAtk [23] leveraged the intrinsic classification capabilities of DMs, facilitating the creation of adversarial prompts. In this work, we treat machine unlearning for DMs as a defensive challenge. Our approach involves fine-tuning the target model to not only unlearn specific concepts but also to enhance its robustness against adversarial prompt attacks.
@@
Adversarial training (AT). In the realm of image classification, adversarial attacks that generate subtle perturbations to fool machine learning (ML) models have long posed a robustness challenge for vision systems [71, 72, 73, 74, 75, 76]. In response, AT (adversarial training) [29], the cornerstone of training-based defenses, conceptualizes defense as a two-player game between the attacker and defender [29, 31, 32, 71, 77, 78, 79, 80, 81, 82, 83, 84, 76]. Additionally, TRADES [31] was proposed to strike a better balance between accuracy and robustness. Further studies [32, 33, 80, 85, 86] demonstrated that unlabeled data and self-training have proven effective in enhancing robustness and generalization in adversarial contexts. To improve the efficiency of AT, past research also proposed adopting more efficient attack methods or fewer steps to generate adversarial examples [71, 87, 88, 89, 90, 91, 92, 93, 94]. In particular, the fast gradient sign method (FGSM) was utilized for adversarial generation in AT [71, 87]. And the gradient alignment strategy was proposed to improve the quality of fast AT [89].
@@ 3 Preliminaries and Problem Statement
Throughout the work, we focus on latent diffusion models (LDMs) [7, 95], which have exceled in text-to-image generation by integrating text prompts (such as text-based image descriptions) into image embeddings to guide the generation process. In LDMs, the diffusion process initiates with a random noise sample drawn from the standard Gaussian distribution . This sample undergoes a progressive transformation through a series of time steps in a gradual denoising process, ultimately resulting in the creation of a clean image . At each time step , the diffusion model utilizes a noise estimator , parameterized by and conditioned on an input prompt (i.e., associated with a textual concept). The diffusion process operates on the latent representation of the image at each time (). The training objective for is to minimize the denoising error as below:
@@
Concept erasure in DMs. DMs, despite their high capability, may generate unsafe content or disclose sensitive information when given inappropriate text prompts. For example, the I2P dataset [10] compiles numerous inappropriate prompts capable of leading DMs to generate NSFW content. To mitigate the generation of harmful or sensitive content, a range of studies [19, 20, 21, 22, 27] have explored the technique of concept erasing or machine unlearning within DMs, aiming to enhance the DM training process by mitigating the impact of undesired textual concepts on image generation.
A widely recognized concept erasing approach is ESD [19], notable for its state-of-the-art (SOTA) balance between unlearning effectiveness and model utility preservation [52]. Unless specified otherwise, we will adopt the objective of ESD for implementing concept erasure. ESD facilitates the fine-tuning process of DMs by guiding outputs away from a specific concept targeted for erasure. Let denote the concept to erase, then the diffusion process of ESD is modified to
where denotes the concept-erased DM, is the originally pre-trained DM, and represents unconditional generation of the model by considering text prompt as empty. Compared to the standard conditional DM [96] (with classifier-free guidance), the second term encourages the adjustment of the data distribution (with erasing guidance parameter ) to minimize the likelihood of generating an image that could be labeled as . To optimize , ESD performs the following model fine-tuning based on (3):
@@
Adversarial prompts against concept-erased DMs. Although concept erasing enhances safety, recent studies [23, 24, 25, 26] have also shown that concept-erased DMs often lack robustness when confronted with adversarial prompt attacks; see Fig. 1 for examples. Let represent a perturbed text prompt corresponding to , obtained through token manipulation in the text space [23, 24] or in the token embedding space [26]. The generation of adversarial prompts can be solved as [23, 24]:
@@
AdvUnlearn: A defensive unlearning setup via AT. The lack of adversarial robustness in concept-erased DMs motivates us to devise a solution that enhances their robustness in the face of adversarial prompts. AT [29] offers a principled algorithmic framework for addressing this challenge. It formulates robust concept erasure as a two-player game involving the defender (i.e., the unlearner for concept erasing) and the attacker (i.e., the adversarial prompt). The original AT constrains the attacker’s objective to precisely oppose the defender’s objective. To loosen this constraint, we consider a generalized AT formulation based on bi-level optimization [92, 97, 98, 99], where the defender and attacker are delineated through the upper-level and lower-level optimization problems, respectively: